mrs.burgos wants to buy at least 30 kilos of pork and beef but can spend no more than 1200. A kilo of pork is 180 and 220 for beef.
1. what mathematical statement represent the situation
2. draw and describe the graph
3. determine the amount of pork and beef mrs burgos needs to buy

Answer by ikleyn(52864)

.
Let X = kilos of pork, Y = kilos of beef.


Inequalities  (= mathematical model)

    x + Y >= 30,

    180X + 220Y <= 1200

    x >= 0,  Y >= y.


The plots are below:





Plots  X + Y = 30 (red),  180X + 220Y = 1200 (green).



The solution domain is the area in QI over the red line and below green line.


It is the EMPTY set (empty intersection). Which means that the problem  HAS NO  solutions.


The set of solutions to the given problem is  EMPTY.